Find the standard deviation of Pandoi vallabie 5. A software package consists of 9 programs, 4 of which must be upgraded. 3 programs are randomly chosen for testing. Construct the distribution law for the number of programs that must be upgraded among chosen.
Answer #1
n = 12, r = 3, p = 4/9, q = 1 - (4/9) = 5/9
x P(x) = C(n, x) p^x q^(n - x)
0 0.171
1 0.412
2 0.329
3 0.088